Given the bulk modulus of sea water is 2.2*109 Pa and the density of surface sea water is 1.025*103 kg/m3. Calculate the volume of 1.025*103 kg of sea water at a depth near 5 km where the pressure is 5.0*107 Pa.
1
Expert's answer
2019-01-30T15:44:08-0500

We have

ΔV/V=ΔP/B

(V-V^')/V=ΔP/B

V^'=V(1-ΔP/B)=m/ρ (1-ΔP/B)

V^'=(1.025∙〖10〗^3)/(1.025∙〖10〗^3 ) (1-(5.0∙〖10〗^7)/(2.2∙〖10〗^9 ))=0.98 m^3.